Balcom Enterprises is planning to introduce a new product that will sell for $145 a unit. Manufacturing cost estimates for 25,000 units for the first year of production are: Direct materials $1,500,000. Direct labor $900,000 (based on $18 per hour × 50,000 hours). Although overhead has not be estimated for the new product, monthly data for Balcom's total production for the last two years has been analyzed using simple linear regression.

The analysis results are as follows: Dependent variable Factory overhead costs Independent variable Direct labor hours Intercept $ 130,000 Coefficient on independent variable $ 5.00 Coefficient of correlation 0.941 R2 0.834 Based on this information, how much is the variable manufacturing cost per unit, using the variable overhead estimated by the regression (assuming that direct materials and direct labor are variable costs)?

Multiple Choice

$106

$83

$88

$97
Answer #1
Direct labor hours per unit 2 =50000/25000
Direct materials per unit 60 =1500000/25000
Direct labor per unit 36 =900000/25000
Variable overhead per unit 10 =5*2
Variable manufacturing cost per unit 106
Option A $106 is correct